# SliM-LLM: Salience-Driven Mixed-Precision Quantization for Large Language Models

### v1

@@ 3.2 Salience-Determined Bit Allocation
Figure 3: Salience in LLaMA-7B.
In the following, we reveal phenomenon of spatial clustering in the distribution of weight salience, which inspires our proposed concept of structured mixed-precision quantization for LLMs, and then present the Salience-Determined Bit Allocation (SBA) technique to optimize bit-width allocation.
3.2.1 Spatial Clustering of Global Salience
We first conduct an empirical investigation into the weight salience distribution. The results reveal that certain channels exhibit higher salience and show tendencies for spatial clustering. As illustrated in Fig. 3, salient clustering are identified around the $2100^{th}$, $3218^{th}$ and $3853^{rd}$ channels within the $2^{nd}$ layer’s attention projection of the LLaMA-7B model. A similar structured pattern is observed near the $600^{th}$, $2200^{th}$ and $3992^{nd}$ channels in the $10^{th}$ layer. Also, clustered salience is detected in other layers (as shown in Fig. 3). More examples of spatial clustering of salience are provided in Appendix F.
Then, we analyze the underlying causes of this phenomenon from a theoretical perspective. According to Definition 1, the salience of weights in LLMs can be quantified numerically by the weight magnitudes and the diagonal elements of the Hessian matrix, which is further approximated by the product of input activations $\boldsymbol{x}\boldsymbol{x}^{\top}$. In LLMs, activations exhibit more extreme outliers, while the numerical differences in weights are relatively slight [44, 32]. Therefore, we propose an analysis of how outlier values in input activations influence the distribution of weight salience:
@@ 3.2.1 Spatial Clustering of Global Salience
Given the input calibration activation $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with an outlier value $x_{p,q}^{*}\gg x_{i,j},\forall i\in[0,t],i\neq p,\forall j\in[0,m],j\neq q$ at the position of token-$p$ and channel-$q$. The diagonal elements of $\boldsymbol{x}\boldsymbol{x}^{\top}$ also shows outlier value at $[\boldsymbol{x}\boldsymbol{x}^{\top}]_{p,p}$, as ${x_{p,q}^{*}}^{2}$ produced by $[\boldsymbol{x}\boldsymbol{x}^{\top}]_{i,i}=\sum_{k=1}^{n}x_{i,k}^{2}$ only appearing at position $(p,p)$, which further leads to the parameter salience larger at the $p^{th}$ channel of weight, where $\delta_{:,p}>\delta_{:,k},\delta_{:,k}=\frac{w_{:,k}^{2}}{[\boldsymbol{H}^{-1}]_{k,k}^{2}},\forall k\in[0,t],k\neq p$.
Theorem 1 elucidates the impact of outlier tokens on the channel-wise distribution of weight salience (detailed proof is provided in Appendix F.1). Additionally, recent studies [32, 45] indicate that outlier tokens in LLMs activations tend to cluster regionally at specific locations, resulting in sequences of consecutive significant tokens. According to Theorem 1, these consecutive tokens will lead to channel clustering results of weight salience, as evidenced by the consecutive salient channels shown in Fig. 3. This also means that when we divide the weight into multiple groups with continuous channels, the overall salience of different groups is different.
@@
Due to the unstructured mixed-precision, which involves additional storage requirements and inference operations, the computational format is not deployment-friendly. However, the spatial clustering of weight salience observed in this section strongly inspired the development of structured mixed-precision strategies with flexible bit-widths while maintaining inference efficiency. Therefore, we aim to allocate bit-width structurally based on group-wise salience differences. This approach not only enhances quantization accuracy but also ensures the deployment efficiency of LLMs.
3.2.2 Salience-Determined Bit Allocation for Structured Group
@@
To allocate optimal bit-widths to each group, we introduce a Salience-Determined Bit Allocation (SBA) technique for mixed-precision LLMs, as depicted in Fig. 1. This technique, predicated on the differences in group salience, determines the optimal bit-width allocation for different groups by minimizing the distance of information entropy with the original weight output.
Specifically, we first utilize the average salience as the importance indicator for each weight group and rank them accordingly. The proposed SBA optimizes the following formula to determine the optimal number of salient-unsalient quantization groups of LLMs:
@@
$\text{Objective}:&\mathop{\mathrm{argmin}}\limits_{g_{1},...g_{n}}\mathcal{D}_{kl}~{}(\boldsymbol{x}\boldsymbol{w}_{f}^{\top}||\boldsymbol{x}\mathcal{Q}(\boldsymbol{w}_{f}|[g_{1},...g_{n}])^{\top}),\\ \text{Constrain}:&|\mathcal{G}_{N-1}|=|\mathcal{G}_{N+1}|,~{}\mathcal{G}_{N-1}=\{g_{i}|g_{i}=N-1\},~{}\mathcal{G}_{N+1}=\{g_{j}|g_{j}=N+1\},\end{aligned}$ (4)
@@
where $\mathcal{D}_{kl}(\cdot||\cdot)$ denotes the Kullback-Leibler (KL) divergence between two outputs, $N$ is the targeted average bit-width. $\mathcal{Q}$ generally represents the quantization and de-quantization processes, employing group-wise mixed precision designated as $[g_{1},...g_{k}]$, where $g_{i}$ represents the bit-width for the $i^{th}$ group and $\mathcal{G}$ is a set of groups with the same bit-width. We apply a compensation constraints strategy to maintain a consistent average bit-width for our SBA. For example, in 2-bit quantization, the groups with the highest salience are quantized to 3-bit. To offset the additional bits, we quantize an equal number of groups with the lowest salience to 1-bit, while the remaining groups are set to 2-bit.
@@
We utilize an effective double-pointer search (more detailed examples in Appendix C) to optimize our objective in Eq. (4). When the weight output channel size is $m$ and group size is $128$, $k=\frac{m}{128}$, the search region for weight is limited to $[0,\frac{k}{2}]$, which is highly efficient with limited searching space, e.g., only 16 iterations are needed in LLaMA-7B. We also provide detailed searching error examples in Appendix C. Notably, SBA diverges from traditional quantization with mean squared error (MSE) in Eq. (3) by instead utilizing the KL divergence as its measure of loss. Beyond simply reducing numerical quantization errors, SBA leverages relative entropy as a mixed bit-width metric, aiming to maximize the mutual information [35] between the quantized and original weights of the LLMs. This approach enhances the model’s capacity for information representation under lower bit quantization.
@@ 3.3.1 Discrete Distribution of Local Salience
Figure 4: Local salience distribution of the $10^{th}$ MHA output layer in LLaMA-7B.
In the aforementioned section, we allocate the bit-width for each group based on the global salience. However, the salience among different elements within the same group still locally exhibits an unstructured difference. Specifically, as depicted in the salience distribution in Fig. 4, within the $10^{th}$ attention output layer of LLaMA-7b, a subset of sparse weights within the comparatively less salient Group-2 still maintains a high level of importance. In LLMs, a small number of weights with outliers may affect the distribution of salience in an unstructured manner, as described in Definition 1. These discrete weights typically account for only approximately 1% of the total weights within the group but play a crucial role in the modeling capability of LLMs.
The existing vanilla quantizers still face the challenge in representing significant weight information, for only considering mean error of all elements within a group. When quantizing weights according to Eq. (1) in group-wise format, a large number of relatively non-salient weights at the intra-group statistical level tend to dominate the parameters generated by the quantizer. This leads to a degradation of salient information within the group, thereby affecting the model performance of LLMs.
3.3.2 Salience-Weighted Quantizer Calibration for Local Salience Awareness
To prevent the degradation of local salient weight information in each group, we propose the Salience-Weighted Quantizer Calibration (SQC), which enhances the expression of salient weights through locally unstructured salience awareness, thereby reducing the quantization error of these significant elements and improving the compressed performance of LLMs. SQC first introduces the calibration parameter $\gamma$ to the quantizer, liberating the perception interval during quantization:
$\Delta=\frac{\gamma(w_{\mathrm{max}}-w_{\mathrm{min}})}{2^{N}-1},~{}z=-\lfloor\frac{\gamma w_{\mathrm{min}}}{\Delta}\rceil$ (5)
where $\gamma$ expands the solution space of the quantizer, subsequently minimizes the loss of quantization in the salience-weighted objective:
$\mathop{\mathrm{argmin}}\limits_{\gamma}~{}\mathcal{L}_{s}(\boldsymbol{w}_{s},\operatorname{dequant}(\hat{\boldsymbol{w}}_{sq},\Delta,z))+\mathcal{L}_{u}(\boldsymbol{w}_{u},\operatorname{dequant}(\hat{\boldsymbol{w}}_{uq},\Delta,z))$ (6)
[…]
SQC effectively mitigates the degradation of local salient weights within groups (more evidence is provided in Appendix D). Additionally, SQC only requires the optimization of $\gamma$ during quantization and without distinguishing unstructured parts in storage and inference stages, thereby avoiding hardware overhead. Combined with SBA, they jointly enhance the awareness of local and global salient weights, capturing significant information in LLMs to improve quantization performance.
Table 1: Quantization results of LLaMA family with statistic quantizer. We report the WikiText2 perplexity in this table, C4 results are shown in Appendix G.
@@ 3.4 Implementation Pipeline of SliM-LLM
We integrate our mixed-precision framework into advanced PTQ methods, such as GPTQ [16] and OmniQuant [38], all of which are deployment-friendly with group-wise quantization. We primarily integrate SBA and SQC into GPTQ to get SliM-LLM. For SliM-LLM+, the SBA is plugged into OmniQuant with a learnable quantizer. The complete pipeline of SliM-LLM is provided in Algorithm 1, and a more detailed implementation is listed in Appendix B.1.
Algorithm 1 Main Framework of SliM-LLM.
Input: $\boldsymbol{w}\in\mathbb{R}^{n\times m}$ FP16 weight
          $\boldsymbol{x}_{F}\in\mathbb{R}^{t\times m}$ calibration data
          $\beta$ group size
          $\lambda$ hessian regularizer
          $N$ average bit-width
Output: $\hat{\boldsymbol{w}}_{q}$ quantized weight
3:  $\hat{\boldsymbol{w}}_{q}\coloneqq 0^{n\times m}$
5:  for $b=0,\beta,2\beta,...$ do
6:     $\boldsymbol{w}^{b}\coloneqq\boldsymbol{w}_{:,b:b+\beta}$
7:     $g_{b}\coloneqq\mathcal{G}[b]$
10:     GPTQ-error compensation:
13:  end for
14:  return $\hat{\boldsymbol{w}}_{q}$
4 Experiments
We evaluated SliM-LLM and SliM-LLM+ under weight-only conditions, focusing on 2/3-bit precisions. Per-channel group quantization is utilized in our framework with 128 set as group size in experiments. Since no back-propagation in SliM-LLM, the quantization is carried out on a single NVIDIA A800 GPU. For SliM-LLM+, we employ the AdamW optimizer, following OmniQuant [38], which is also feasible on a single A800. We randomly select 128 samples from WikiText2 [30] as calibration data, each with 2048 tokens.
Models and Evaluation. To comprehensively demonstrate the low-bit performance advantages of SliM-LLM and SliM-LLM+, we conduct experiments across OPT [49], LLaMA [41], LLaMA-2 [42] and LLaMA-31. We employ the perplexity as our evaluation metric, which is widely recognized as a stable measure of language generation capabilities [16, 26, 20, 37, 38, 6, 13, 21], particularly in compression scenarios. Experiments are carried out on the WikiText2 [30] and C4 [36]datasets. Furthermore, to assess the practical application capabilities of quantized LLMs, we also evaluate their accuracy on zero-shot benchmarks, including PIQA [2], ARC [10], BoolQ [9], and HellaSwag [10].
Baseline. Since SliM-LLM and SliM-LLM+ are efficient PTQ approaches without additional training or fine-tuning, QAT and re-training methods are not within the comparison range of our work. The experiments evaluate existing advanced quantization methods and GPU-friendly computations, including vanilla round-to-nearest (RTN), GPTQ [16], AWQ [26]. And mixed-precision quantization techniques, including PB-LLM [37] ($\frac{1}{7}\times 8$-bit+$\frac{6}{7}\times 1$-bit), LLM-MQ [24], and APTQ [18], as well as the codebook-based method QuIP [6] are also compared in this work. We compare SliM-LLM+ with gradient optimizer-based methods such as OmniQuant [38] and AffineQuant [28].
4.1 Main Results.
We show experiments within the LLaMA family in this section and detailed results for the OPT models are available in Appendix G. For language generation tasks, as depicted in Tab. 1, SliM-LLM markedly outperforms its backbone GPTQ, particularly under the 2-bit. Specifically, on LLaMA-7B, SliM-LLM achieves a 90% decrease in perplexity, while on LLaMA-3-8B, it improves performance by 81%. In comparison with the unstructured mixed-precision PB-LLM and the codebook-based QuIP method, SliM-LLM further reduces the perplexity by 41%~51%. As shown in Tab. 1, the performance of SliM-LLM+ is still ahead compared to OmniQuant and AffineQuant, further proving the effectiveness and of the mixed-precision framework proposed in our work. We also provide dialogue examples of 2-bit instruction fine-tuning Vicuna-13B [8] and LLaMA-13B in Appeandix H.
Morever, our method exhibits zero-shot advantages at 2-bit, as shown in Tab. 3, where SliM-LLM and SliM-LLM+ still outperforms other methods. For instance, compared with GPTQ and OmniQuant, our approach achieves an average improvement of 4.19% and 1.91% on LLaMA-7B. Meanwhile, for LLaMA-65B, 2-bit SliM-LLM and SliM-LLM+ is close to FP16 results (less than 6% degradaion in accuracy). Overall, our proposed mixed-precision framwork demonstrates superior performance across different model sizes, with its advantages becoming increasingly significant at lower bit-width.
Table 2: Quantization results of LLaMA-1 and LLaMA-2 models with learnable quantizer. We report the WikiText2 perplexity in this Table, C4 results are shown in Appendix G.

### v2

@@ 3.2 Salience-Determined Bit Allocation
We reveal the spatial clustering of weight salience, which inspires our proposed concept of group-wise mixed-precision quantization for LLMs, and then introduce the Salience-Determined Bit Allocation (SBA) technique to allocate the optimal precision to each group.
3.2.1 Spatial Distribution of Global Salience
Figure 3: Salience weight distribution in layer-2 and layer-10 of LLaMA-7B.
We first conduct an empirical investigation into the weight salience distribution. The results reveal that certain channels exhibit higher salience and show tendencies to spatial clustering. As illustrated in Fig. 3, salient clustering is identified around the $2100^{th}$, $3218^{th}$ and $3853^{rd}$ channels within the $2^{nd}$ layer’s attention projection of the LLaMA-7B model. A similar structured pattern is observed in other layers. Also, clustered salience is detected in other layers (as shown in Fig. 3). More examples are provided in Appendix G.
Then, we analyze the underlying reasons for this phenomenon. According to Definition 3.1,the salience of weights is proportional to the magnitude of the weights and the trace of the Hessian matrix, which can be approximated by the product of input activations $\boldsymbol{x}^{\top}\boldsymbol{x}$. In LLMs, activations exhibit extreme outlier channels, while the numerical differences in weights are relatively slight (Xiao et al., 2023a; Nrusimha et al., 2024). Therefore, we propose an analysis of how the outlier channels in activations influence the distribution of weight salience:
@@ 3.2.1 Spatial Distribution of Global Salience
Given the input calibration activation $\boldsymbol{x}\in\mathbb{R}^{t\times m}$ with an outlier channel $\boldsymbol{x_{:,p}^{*}}\gg\boldsymbol{x}_{:,j},\forall j\in[0,m],j\neq p$ at the position of channel-$p$. The trace elements of $\boldsymbol{H}=\boldsymbol{x}^{\top}\boldsymbol{x}$ will show great outlier value at $(p,p)$, where $\boldsymbol{H}_{p,p}\gg\boldsymbol{H}_{j,j},\forall j\in[0,m],j\neq p$, as $\boldsymbol{H}_{p,p}$ is produced by $[\boldsymbol{x_{:,p}^{*\top}}\boldsymbol{x_{:,p}^{*}}]=\sum_{i=0}^{t}x_{i,p}^{*2}$, which further leads to the parameter salience larger at the $p^{th}$ channel of weight, where $\delta_{:,p}>\delta_{:,k},\delta_{:,k}=\frac{w_{:,k}^{2}}{[\boldsymbol{H}^{-1}]_{k,k}^{2}},\forall k\in[0,t],k\neq p$.
Theorem 1 elucidates the influence of outlier activation on the distribution of channel salient weights (detailed proof in the Appendix G.1). Furthermore, recent research indicates that outlier channels in LLMs activations consistently appear in fixed yet clustered patterns (Nrusimha et al., 2024). According to Theorem 1, these consistently occurring anomalous activations result in the distribution of salient weights, as depicted in Fig. 3. Then, during group-wise quantization, the average salience of each group shows different features.
@@
Meanwhile, previous unstructured mixed-precision, incurred additional storage requirements and computational overheads, affecting the real-time inference. However, the strong spatial structured characteristics observed in the salient of weights in this section strongly inspire us to first develop a group-wise mixed-precision strategy within the weight matrix while maintaining inference efficiency. Therefore, we aim to allocate bit-widths based on intra-group salient disparities, which not only enhances quantization accuracy but also ensures the inference efficiency of LLMs with structured bit-widths saving and dequantization.
3.2.2 Group-Wise Bit Allocation
@@
To allocate optimal bit-widths to each group, we introduce a Salience-Determined Bit Allocation (SBA) technique for mixed-precision LLMs, as depicted in Fig. 2. This technique, predicated on the differences in group salience, determines the optimal bit-widths allocation for different groups by minimizing the distance of information entropy with the original weight output. Specifically, we first utilize the average salience as the importance indicator for each weight group and rank them accordingly. Then, the proposed SBA optimizes the bit allocation following:
@@
$\text{Objective}:&\mathop{\mathrm{argmin}}\mathcal{D}_{kl}~{}(\boldsymbol{x}\boldsymbol{w}_{f}^{\top}~{}||~{}\boldsymbol{x}(\hat{\boldsymbol{w}}_{sba})^{\top}),\\ &\mathrm{where}~{}\hat{\boldsymbol{w}}_{sba}=[\hat{\boldsymbol{w}}_{0,b_{0}},\hat{\boldsymbol{w}}_{1,b_{1}}...\hat{\boldsymbol{w}}_{k-1,b_{k-1}},\hat{\boldsymbol{w}}_{k,b_{k}}]\\ \text{Constrain}:&|\mathcal{G}_{N-1}|=|\mathcal{G}_{N+1}|,\\ &\mathrm{where}~{}\mathcal{G}_{N-1}=\{b_{i}|b_{i}=N-1\},\\ &\mathcal{G}_{N+1}=\{b_{j}|b_{j}=N+1\}\end{aligned}$ (4)
@@
where $\mathcal{D}_{kl}(\cdot||\cdot)$ denotes the Kullback-Leibler (KL) divergence between two outputs, $\hat{\boldsymbol{w}}_{f}^{sba}$ generally represents the de-quantization results of weight, employing group-wise mixed-precision designated as $[\hat{\boldsymbol{w}}_{0,b_{0}},\hat{\boldsymbol{w}}_{1,b_{1}}...\hat{\boldsymbol{w}}_{k-1,b_{k-1}},\hat{\boldsymbol{w}}_{k,b_{k}}]$, where $b_{i}$ represents the bit-widths for the $i^{th}$ group and $\mathcal{G}$ is a set of groups with the same bit-width, $N$ is the targeted average bit-width. We apply a compensation constraints strategy to maintain a consistent average bits for our SBA. For example, in 2-bit quantization, the more salient groups are quantized to 3-bit. To offset the additional bits, we quantize an equal number of groups with the lower salience to 1-bit ($|\mathcal{G}_{N-1}|=|\mathcal{G}_{N+1}|$), while the remaining groups are set to 2-bit.
@@
We utilize an effective double-pointer search (more detailed examples in Appendix C) to optimize our objective in Eq. (4). When the weight output channel size is $m$ and group size is $128$, $k=\frac{m}{128}$, the search region for weight is limited to $[0,\frac{k}{2}]$, which is highly efficient with limited searching space, e.g., only 16 iterations are needed in LLaMA-7B. We also provide detailed searching error examples in Appendix C. Notably, SBA diverges from traditional quantization with mean squared error (MSE) in Eq. (3) by instead utilizing the KL divergence as its measure of loss. Compared to using mean squared error (MSE) for weights, SBA utilizes the KL divergence of block outputs as a precision allocation metric, aiming to align the distribution of the LLM’s output activation matrix with that of the quantized activation. This approach improves the model’s information representation under low-bit quantization, enabling more effective bit-widths allocation. While HAWQ v2 (Dong et al., 2019) employs Integer Linear Programming (ILP) to allocate bit-widths for layers, its method can be adapted to group-wise targets. Detailed comparisons between SBA and ILP are provided in Section 4.2.
@@ 3.3.1 Discrete Distribution of Local Salience
Figure 4: Local salience distribution of the $10^{th}$ MHA output layer.
In the aforementioned section, we group-wisely allocate the bit-widths for each group based on the global salience. To maintain the efficiency of quantized inference, we employ a commonly used sequential structured grouping (Frantar et al., 2022; Lin et al., 2023; Shao et al., 2023). However, this group-wise mixed-precision also leads to differences in salience among the various elements within the same group. Specifically, as the salience distribution in Fig. 4, within the $10^{th}$ attention output layer of LLaMA-7B, a subset of sparse weights within the comparatively less salient Group-2 (Fig. 4) still maintains a high level of importance. In LLMs, a small number of weight elements with outliers affect the local distribution of salience. These discrete weights typically account for only approximately 1% within the group but play a crucial role in the modeling capability of LLMs.
Table 1: Quantization results of LLaMA family with statistic quantizer. We report the WikiText2 perplexity in this table, C4 results are shown in Appendix H. ‘-’ denotes that the selected works did not give the results on listed models or the codes.
@@
Vanilla quantizers struggle to represent local salient information as they only consider the mean error of all elements within a group. During group-wise quantization (Eq. 1), non-salient weights often dominate, degrading critical information and negatively impacting LLM performance.
3.3.2 Salience-Weighted Quantizer
To prevent the degradation of local salient information in each group, we propose the Salience-Weighted Quantizer Calibration (SQC), which enhances the expression of salient weights through local salience awareness, thereby reducing the quantization error of these significant elements and improving the compressed performance of LLMs.
Based on a common observation (Dettmers et al., 2023; Huang et al., 2024a), the proportion of relatively salient weights in each group is only 1-5%. Therefore, we employ the 3-$\sigma$ rule for a mask to select the salience part ($\boldsymbol{w}<(\mu-3\sigma)\cup\boldsymbol{w}>(\mu+3\sigma)$) in each group (Fig. 2), which accounts for about 1% elements. After the selection, we get $\boldsymbol{w}_{i}=\boldsymbol{w}_{i}^{s}\cup\boldsymbol{w}_{i}^{us}$, where $\boldsymbol{w}_{i}^{s}$ is the salient part and $\boldsymbol{w}_{i}^{us}$ represents the non-salient elements within group $i$. To effectively keep the information of local salient weights, SQC first introduces the calibration parameter $\tau$ to the SQC quantizer, liberating the perception interval during quantization. Then we define the local salience awareness loss of the SQC quantizer through calibration:
$\mathop{\mathrm{argmin}}\limits_{\tau}~{}||\boldsymbol{w}_{i}^{s}-\tau\cdot s\{\mathcal{Q}(\boldsymbol{w}_{i}^{s},\tau\cdot s,\tau\cdot z)-\tau\cdot z\}||_{2}^{2}+$ (5)
$||\boldsymbol{w}_{i}^{us}-\tau\cdot s(\mathcal{Q}(\boldsymbol{w}_{i}^{us},\tau\cdot s,\tau\cdot z)-\tau\cdot z)||_{2}^{2}$
[…]
Compared to traditional quantizer calibration methods, SQC effectively mitigates the degradation of intra-group local salient weights caused by general average loss by enhancing the loss sensitivity to salient elements during the calibration (more experiments are detailed in Appendix E). Moreover, the SQC process allows $\boldsymbol{w}_{i}^{s}$ and $\boldsymbol{w}_{i}^{us}$ to share a set of parameters $\tau^{*}s$ and $\tau^{*}z$, eliminating the need to differentiate intra-group weights during storage and inference. This facilitates straightforward group-wise dequantization calculations, thereby avoiding the hardware overhead associated with element-wise bitmap and unstructured grouping. SQC and SBA capture both local salient weight information within groups and global weight combinations across groups, effectively preserving critical information and maintaining LLM performance at extremely low bit-widths.
@@ 3.4 Implementation Pipeline of SliM-LLM
We integrate our mixed-precision framework into advanced PTQ methods, such as GPTQ (Frantar et al., 2022) and OmniQuant (Shao et al., 2023), all of which are inference-friendly with group-wise quantization. We primarily integrate SBA and SQC into GPTQ to get SliM-LLM. For SliM-LLM+, the SBA is plugged into OmniQuant with a learnable quantizer. The plugging pipeline of SliM-LLM is provided in Algorithm 1 (line 4 and line 9), detailed functions are shown in Algorithm 2.
Table 2: Quantization results of LLaMA-1 and LLaMA-2 models with learnable quantizer. We report the WikiText2 perplexity in this Table, C4 results are shown in Appendix H.
@@
4 Experiments
We evaluated SliM-LLM and SliM-LLM+ under weight-only conditions, focusing on 2/3-bit precisions. Per-channel group quantization is utilized in our framework with $groupsize=128$ in experiments. Since no back-propagation in SliM-LLM, the quantization is carried out on a single NVIDIA A800 GPU. For SliM-LLM+, we employ the AdamW optimizer, following OmniQuant (Shao et al., 2023), which is also feasible on a single A800. We randomly select 128 samples from WikiText2 (Merity et al., 2016) as calibration data, each with 2048 tokens.
Models and Evaluation. To comprehensively demonstrate the low-bit performance advantages of SliM-LLM and SliM-LLM+, we conduct experiments across OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b) and LLaMA-3. We employ the perplexity as our evaluation metric, which is widely recognized as a stable measure of language generation capabilities (Frantar et al., 2022; Lin et al., 2023; Huang et al., 2024a; Shang et al., 2023; Shao et al., 2023; Chee et al., 2024; Egiazarian et al., 2024; Huang et al., 2024b), particularly in compression scenarios. Experiments are carried out on the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020)datasets. Furthermore, to assess the practical application capabilities of quantized LLMs, we also evaluate on zero-shot benchmarks.
4.1 Main Results
We show experiments within the LLaMA family in this section and results for the OPT models are available in Appendix H. For language generation tasks, as depicted in Tab. 1, SliM-LLM markedly outperforms its backbone GPTQ, particularly under the 2-bit. Specifically, on LLaMA-7B, SliM-LLM achieves a 90% decrease in perplexity, while on LLaMA-3-8B, it improves by 81%. In comparison with the element-wise mixed-precision PB-LLM and the codebook-based QuIP method, SliM-LLM further reduces the perplexity by 41%~51%. As shown in Tab. 1, the performance of SliM-LLM+ is still ahead compared to OmniQuant and AffineQuant. We also provide dialogue examples of 2-bit instruction fine-tuning Vicuna-13B (Chiang et al., 2023) and LLaMA-13B in Appeandix I. Our method demonstrates zero-shot advantages at 2-bit, as shown in Tab. 3, where SliM-LLM and SliM-LLM+ outperform other methods. For example, compared to GPTQ and OmniQuant, our approach improves LLaMA-7B performance by 4.19% and 1.91% on average. On LLaMA-65B, 2-bit SliM-LLM and SliM-LLM+ achieve accuracy within 6% of FP16. To further showcase the general performance of SliM-LLM, we also compare the low-bit results on multi-modal models (Tab. 15), where SLiM-LLM presents leading accuracy on 4 benchmarks.